OpenStudy (anonymous):

Solve for x: 4/3 x + 2/3 = 6 x = 4 x = 3 x = 1 x = 0

OpenStudy (mathstudent55):

First, multiply the entire equation by 3. This will get rid of the denominators.

OpenStudy (anonymous):

(4/3)x=6-2/3 (4/3)x=16/3 x=(16/3)(3/4) x=16/4 x=4

OpenStudy (anonymous):

3 = x

OpenStudy (anonymous):

eyh, yea, my bad -_-

OpenStudy (anonymous):

wait, no, x=4

OpenStudy (anonymous):

Thank You :)

OpenStudy (anonymous):

no problem :D